# Large Language Models are Easily Confused: A Quantitative Metric, Security Implications and Typological Analysis

### v1

@@ 4.4 Discussion
In response to RQ1, we proposed an effective metric Language Confusion Entropy, through which we identified several patterns contributing to language confusion. These include prompt complexity, imbalanced distributions of training sources, and similarities within language families — all play a significant role in language confusion. Furthermore, our findings indicate that these factors strongly correlate with inversion performance and the pretraining languages in LLMs.
@@
In response to RQ2, our findings confirm the hypothesis that language confusion is an inherent vulnerability of LLMs, and can be influenced and explained by linguistic typological variation. Thus, there is a potential to leverage language similarities as a prior for LLM alignment and security. For instance, our findings suggest that language confusion in LLMs could be reduced by incorporating languages with distinct semantic typological features. As observed in Table 3, ensuring that the model is exposed to languages with greater differences in colexification patterns may enhance its ability to distinguish between them. This strategy could promote more resilient LLMs, as we have shown that models are less likely to confuse typologically dissimilar languages. Hence, exploring typology-aware design strategies could provide both offensive and defensive insights in LLM security.
@@ 5 Conclusion and Future Work
Addressing the challenge of language confusion, we introduce Language Confusion Entropy, a novel metric that quantifies language confusion by re-weighting language distributions and emphasizing long-tail patterns. This metric captures language confusion in multilingual LLM tasks, revealing patterns of uncertainty in both training and evaluation phases. Our findings show strong correlations between language confusion and semantic similarities among languages, with less confusion observed in low-resource languages and when training incorporates diverse scripts and language families. These insights confirm that language confusion fundamentally impacts LLMs and suggest linguistic typology as a potential tool for enhancing model security. In future work, we aim to apply these findings to practical applications, such as developing typology-aware defense mechanisms and attack strategies to improve LLM alignment and security.
@@
A core limitation of this work is that our analysis and downstream implications can only be carried out on languages that are represented in typological databases. While these databases typically have high coverage, this still leaves some languages in the dark. When this work inspires downstream solutions in terms of defense mechanisms, undocumented languages may not benefit from these advances.
@@
This work adheres to the ACL ethics guidelines. We investigate language confusion and link findings to security vulnerabilities of low-resource languages, including those using non-Latin scripts and with diverse typologies. Our work highlights how these factors can be used to potentially improve the security of low-resource language technology. We encourage the community to incorporate a broader range of languages in NLP security research, to ensure that low-resource languages are also covered by defense mechanisms developed in the future.
@@ Appendix A Datasets for Language Confusion
Langauge Confusion Benchmark
@@
Textual embedding inversion has presented a standing challenge in LLM security, where the private texts can be reconstructed from evasdroped embeddings from Embeddings as a Service (EaaS), by training an attacker model based on the embeddings extracted from the black-box embedders (Song and Raghunathan, 2020; Lyu et al., 2020; Kim et al., 2022; Morris et al., 2023; Chen et al., 2024b, a). However, most work was done in monolingual settings, mostly in English, other than the recent work expands the language space to four Romance and Germanic languages in Latin script  (Chen et al., 2024b) and in Chen et al. (2024a), the inversion atacks are extended to 20 languages across 8 families and 12 scripts (see Table 7). The trained inversion attack model consists of a base model and a corrector model, where a base model is a text-to-text generation model, while a corrector model is used to bring closer the generated embeddings and attacked embeddings in the embedding space. While in the evaluation phase, three stages are reported: base, step1 (corrector model) and step50+sbeam8 (corrector model with beam search with sequence length 8). The inversion model is trained with mT5 (Xue, 2020) as base model and multilingual-e5-base  [^5] Huggingface: intfloat/multilingual-e5-base as black-box encoder. The samples of the curated dataset are shown in Table 8.
@@
Table 7: Languages and their Language Characteristics, i.e., Language Family, Language Script, Directionality of the Script, Number of Training Samples for Inversion Models, Word Order of Subject, Object and Verb in Multilingual Inversion Attack (Chen et al., 2024a) and the Percentage of the language in Pre-training data in mT5 (Xue, 2020).
@@
Table 8: Examples of Dataset MTEI.
@@ Appendix B Language Graphs for Language Similarities
Table 15: KL Divergence between Language Similarity Graphs and the Language Confusion Matrices for Target/ Eval Languages from LCB and Inversion Tasks. The best results (lowest) are bolded, the second best are underlined.

### v2

@@ 5 Discussion
In response to RQ1, we proposed an effective metric Language Confusion Entropy, through which we identified several patterns contributing to language confusion. These include prompt complexity, imbalanced distributions of training sources, and language similarities - all play a significant role in language confusion. Furthermore, our findings indicate that these factors strongly correlate with inversion performance and the pretraining languages in LLMs.
Inherent vulnerabilities in LLMs stem from intrinsic design, training processes, or model architectures, which are not directly caused by attacks or improper use. However, external attacks can amplify and leverage the vulnerabilities, which are directly reflected in embedding inversion attacks. Further, decoding unexpected languages disrupts the user experience. We have discussed that i) language confusion stems from training data inequality, ii) the phenomena is pervasive from related works, iii) language confusion is well-captured by the proposed LCE and shows consistent trends across different architectures of LLMs and datasets and experimental settings, and (iv) can be influenced and explained by linguistic typological variations. These affords a positive response to RQ2.
@@
Moreover, there is a potential to leverage language similarities as a prior for LLM alignment and security. For instance, Table 3 shows that the language similarities from colexification patterns afford a strong correlation with language confusion (with low KL divergence), which indicates that LLMs easily confuse languages that contain words that are crosslingually capturing the same senses. When an LLM is exposed to multilingual data with more distinct colexification patterns, it could enhance its ability to distinguish them and make them more resilient against language confusion. This strategy could promote more resilient LLMs, as we have shown that models are less likely to confuse typologically dissimilar languages. Hence, exploring typology-aware design strategies could provide both offensive and defensive insights in LLM security.
Potential Misuse
The Language confusion metric identifies uncertainty in language identification, highlighting areas where LLMs may be prone to errors. Our findings show that language similarities correlate with language confusion patterns in LLMs. When similar languages lack robust safety measures compared to well-protected high-resource languages, they could be exploited for crosslingual attacks in a more targeted manner, such as backdoors and jailbreaking (Zou et al., 2023; Li et al., 2024). Additionally, the ability of LLMs to switch between languages may pose risks where safety measures aren’t consistently implemented across all languages. The Language confusion metric identifies uncertainty in language identification, highlighting areas where LLMs may be prone to errors. Our findings show that language similarities correlate with language confusion patterns in LLMs. When similar languages lack robust safety measures compared to well-protected high-resource languages, they could be exploited for crosslingual attacks in a more targeted manner, such as jailbreaking (He et al., 2024) and backdoors (Wang et al., 2024). Additionally, the ability of LLMs to switch between languages may pose risks where safety measures aren’t consistently implemented across all languages. It has been demonstrated that LLMs are particularly vulnerable to crosslingual attacks in related work (Wang et al., 2024; He et al., 2024) and more recently (Poppi et al., 2024), mainly because the conventional defense mechanisms designed for monolingual settings are ineffective in multilingual settings. To enhance LLM security, adversarial exploits can be detected by monitoring high language confusion entropy, especially in crosslingual settings. Our work suggests that language confusion quantification and its connection to language similarities can be leveraged to raise awareness of such vulnerabilities, and also provide potential revenue for developing mitigation strategies.
@@ 6 Conclusion and Future Work
Addressing the challenge of language confusion, we introduce Language Confusion Entropy, a novel metric that quantifies language confusion by re-weighting language distributions and emphasizing long-tail patterns. This metric captures language confusion in multilingual LLM tasks, revealing patterns of uncertainty in both training and evaluation phases. Our findings show strong correlations between language confusion and semantic similarities among languages, with less confusion observed in low-resource languages and when training incorporates diverse scripts and language families. These insights confirm that language confusion fundamentally impacts LLMs and suggest linguistic typology as a potential tool for enhancing model security. Detecting language confusion enables smoother and more precise interactions in multilingual contexts while enhancing user trust in LLM-based AI systems across domains such as legal services and healthcare. In future work, we aim to apply these findings to practical applications, such as developing typology-aware defense to improve LLM alignment and security. Key applications include cross-lingual chatbots, translation services, detecting code-switching, and improving multilingual speech recognition by reducing ambiguity.
@@
A core limitation of this work is that some of our analysis and downstream implications can only be carried out on languages that are represented in typological databases. When part of the work, limited to typological databases, inspires downstream solutions in terms of defense mechanisms, undocumented languages may not benefit from these advances. However, we have also increased coverage of languages in constructing language graphs using data-driven methods. Our core method is also not limited to any typological database.
@@
This work adheres to the ACL ethics guidelines. We investigate language confusion and link findings to security vulnerabilities of low-resource languages, including those using non-Latin scripts and with diverse typologies. The potential misuse has been extensively discussed. Our work highlights how these factors can be used to potentially improve the security of low-resource language technology. We encourage the community to incorporate a broader range of languages in NLP security research, to ensure that low-resource languages are also covered by defense mechanisms developed in the future.
Acknowledgements
YC and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454).
@@ Appendix A Datasets for Language Confusion
Language Confusion Benchmark
@@
Textual embedding inversion has presented a standing challenge in LLM security, where the private texts can be reconstructed from evasdroped embeddings from Embeddings as a Service (EaaS), by training an attacker model based on the embeddings extracted from the black-box embedders (Song and Raghunathan, 2020; Lyu et al., 2020; Kim et al., 2022; Morris et al., 2023; Chen et al., 2024, 2025). However, most work was done in monolingual settings, mostly in English, other than the recent work expands the language space to four Romance and Germanic languages in Latin script  (Chen et al., 2024) and in Chen et al. (2025), the inversion atacks are extended to 20 languages across 8 families and 12 scripts (see Table 7). The trained inversion attack model consists of a base model and a corrector model, where a base model is a text-to-text generation model, while a corrector model is used to bring closer the generated embeddings and attacked embeddings in the embedding space. While in the evaluation phase, three stages are reported: base, step1 (corrector model) and step50+sbeam8 (corrector model with beam search with sequence length 8). The inversion model is trained with mT5 (Xue, 2020) as base model and multilingual-e5-base  [^5] Huggingface: intfloat/multilingual-e5-base as black-box encoder. The samples of the curated dataset are shown in Table 8.
@@
Table 7: Languages and their Language Characteristics, i.e., Language Family, Language Script, Directionality of the Script, Number of Training Samples for Inversion Models, Word Order of Subject, Object and Verb in Multilingual Inversion Attack (Chen et al., 2025) and the Percentage of the language in Pre-training data in mT5 (Xue, 2020).
@@
Table 8: Examples of Dataset MTEI. The probabilities for unidentified languages are omitted.